OpenStudy (anonymous):

Find cot θ if csc θ = (square root of 17)/4 and tan θ > 0

OpenStudy (anonymous):



OpenStudy (anonymous):

you need the third side, which you get by pythagoras it is \(\sqrt{\sqrt{17}^2-4^2}=\sqrt{17-16}=1\)

OpenStudy (anonymous):

So, the answer is 1?? Because my answer choices are: sqrt 17 1/4 (sqrt 17)/17 17 thanks for your help btw (: @satellite73

zepdrix (zepdrix):

\[\large \csc \theta = \frac{\sqrt {17}}{4} \qquad \rightarrow \qquad \frac{1}{\sin \theta}= \frac{\sqrt {17}}{4} \qquad \rightarrow \qquad \sin \theta=\frac{4}{\sqrt {17}}\] While Sat has the right idea, I think we have the triangle setup incorrectly :)

zepdrix (zepdrix):

Finding the missing side as Sat did gives you 1

zepdrix (zepdrix):

\[\cot \theta = \frac{adjacent}{opposite}\]

OpenStudy (anonymous):

Ohhhh... okay! So, the answer would be 1/4? (:

zepdrix (zepdrix):

Yay good job! \c:/

OpenStudy (anonymous):

Thank you so much!! :D